y=10x^2 + x
5x+y=4

Answer by checkley71(8403)

Y=4-5X THEN 4-5X=10X~2+X OR 10X~2+6X-4=0 OR 5X~2+3X-2=0 OR (5X-2)(X+1)=0 OR
5X=2 OR X=2/5 & X=-1 THEN
Y=4-5*2.5 OR Y=2 & Y=4-5(-1) OR Y=9